OpenStudy (anonymous):

Find the indefinite integral: (t + 1) / (t^2 + 4) Show steps.

OpenStudy (anonymous):

Let t = 2tan x dt/dx = 2 sec^2 x dt = 2 sec^2 x dx Therefore \begin{eqnarray*} \int{\frac{t+1}{t^2+4}}dt&=&\int{\frac{2\tan x + 1}{4\tan^2 x + 4}2\sec^2 x} dx \\ &=&\int{\frac{2\tan x + 1}{4\sec^2 x}2\sec^2 x}dx \\ &=&\int{\frac{2\tan x + 1}{2}} dx \end{eqnarray*} which should then be easy to integrate.

OpenStudy (anonymous):

Sweet trig, thanks!

OpenStudy (nikvist):

\[\int\frac{t+1}{t^2+4}dt=\int\frac{t}{t^2+4}dt+\int\frac{1}{t^2+4}dt=\]\[=\frac{1}{2}\int\frac{d(t^2+4)}{t^2+4}+\frac{1}{4}\int\frac{1}{(t/2)^2+1}dt=\frac{1}{2}\ln{(t^2+4)}+\frac{1}{2}\arctan\frac{t}{2}+C\]